# latexText page

### v1

@@ Abstract
Significant progress has been achieved on the improvement and downstream usages of the Contrastive Language-Image Pre-training (CLIP) vision-language model, while less attention is paid to the interpretation of CLIP. We propose a Gradient-based visual and textual Explanation method for CLIP (Grad-ECLIP), which interprets the matching result of CLIP for specific input image-text pair. By decomposing the architecture of the encoder and discovering the relationship between the matching similarity and intermediate spatial features, Grad-ECLIP produces effective heat maps that show the influence of image regions or words on the CLIP results. Different from the previous Transformer interpretation methods that focus on the utilization of self-attention maps, which are typically extremely sparse in CLIP, we produce high-quality visual explanations by applying channel and spatial weights on token features. Qualitative and quantitative evaluations verify the effectiveness and superiority of Grad-ECLIP compared with the state-of-the-art methods. Furthermore, a series of analysis are conducted based on our visual and textual explanation results, from which we explore the working mechanism of image-text matching, the strengths and limitations in attribution identification of CLIP, and the relationship between the concreteness/abstractness of a word and its usage in CLIP. Finally, based on the ability of explanation map that indicates text-specific saliency region of input image, we also propose an application with Grad-ECLIP, which is adopted to boost the fine-grained alignment in the CLIP fine-tuning. The code of Grad-ECLIP is available here: https://github.com/Cyang-Zhao/Grad-Eclip.
Index Terms:
gradient-based explanation, visual and textual explanation, explainable AI, contrastive language-image pre-training, fine-grained understanding, open-vocabulary detection, deep learning
1 Introduction
@@
Recently, by learning the representations for matching caption text and its corresponding image, the Contrastive Language-Image Pre-training (CLIP) model [1] has introduced a simple and effective dual-encoder pre-training paradigm for the interaction between natural language processing and computer vision. CLIP significantly improves the performance on various downstream tasks, such as classification [2, 3], retrieval [4] and segmentation [5, 6], with zero-shot and fine-tuning methodologies. Inspired from CLIP, multi-modal pre-training has been further developed by exploring different perspectives, including unifying vision-language understanding and generation [7, 8], prompt design [9, 10], and region-aware enhancement [11, 12, 13]. Although researchers devote many efforts into improving multi-modal pre-training or exploring the usages in downstream tasks, less attention has been focused on the interpretation or explanation of CLIP.
Previous visual explanation works have considered interpreting the transformer architecture used by CLIP. Attention Rollout [14] generates explanations by aggregating attention maps computed along the forward pass of the model. Relevance-based methods [15, 16] apply Layer-wise Relevance Propagation (LRP) [17] and also rely on the attention mechanism in the model architecture. Since Rollout and many LRP variants are class-agnostic, Transformer interpretability [15] and Generic Attention-Model Explainability (GAME) [16] build class-specific relevance-based explanations using the self-attention or co-attention. However, just treating CLIP as a vision transformer (ViT) and generating visual explanations based on self-attention sometimes leads to confusing results because of sparse attention maps (see Fig. 1e-g).
@@
Figure 1: Visual and textual explanations of CLIP for the image with the text “A dog is playing with frisbee” using (a) CLIPSurgery [18]; (b) MaskCLIP [19]; (c) Grad-CAM [20]; (d) RISE [21]; (e) raw attention in the last layer; (f) Rollout [14]; (g) GAME [16]; (h) M2IB [22]; and (i) Our Grad-ECLIP. For (e) to (i), textual explanations on the sentence are shown, where the degree of green color represents the word importance. Other methods (a-d) are not applicable on text.
@@
ECLIP [23] and CLIPSurgery [18] (see Fig. 1a) explores explanations for CLIP by computing an image-text similarity map, and solve the counter-intuitive problem that background patch features get higher similarity with the text feature than the foreground. However, to obtain reasonable similarity maps, new additional projection layers or changing the structure of the original CLIP are required. Although the parameters of CLIP encoders are frozen, learning more black-box parameters with extra data or modifying original model architecture makes the explanation less interpretable. MaskCLIP [19] also provides a technique to calculate class-specific image-text similarity map. By passing the value features of the last attention layer through later linear layers as image patch features, the similarity map is able to localize the concept in the text (see Fig. 1b), but has noisy backgrounds and confusingly highlights points on the locations unrelated to the explained target. The disadvantage of these similarity-map methods is that they are only forward processing, and the attended features are not necessarily used in the final prediction.
@@
To better focus on the discriminative features used in the prediction, gradient-based methods with class-activation maps (CAM) [24], such as Grad-CAM [20], Layer-CAM [25] and FullGrad [26], consider the gradient of the prediction with respect to features from a CNN layer as weights, and locates the class-specific discriminative regions by weighted aggregation of the features maps. Fig. 1c shows the visualization when adapting Grad-CAM on CLIP, where the cosine similarity of image-text pair is adopted as the prediction and the gradients are calculated w.r.t. the patch tokens from the ViT layers. Since there are no gradients w.r.t. the patch tokens in final layer because they are not involved in the calculation of the matching score, feature outputs from the penultimate layer of ViT are adopted. The results of Grad-CAM do not well explain CLIP, and suffer from the same problem of highlighting unrelated points as MaskCLIP, suggesting that the layer features of ViT are unsuitable for CAM methods.
@@
In this paper, we explore a more effective way to interpret CLIP, by analyzing how CLIP obtains the final feature embedding, and deriving the relationship between the image and text embedding similarity score and intermediate features via a series of approximations. Based on the CAM principle, we then propose a novel gradient-based visual explanation method for CLIP (Grad-ECLIP), which generates the importance heat map by aggregating the intermediate image features with result-related channel and spatial importance weights. Our proposed method uses the gradients of the image-text matching score w.r.t. the attention layer as the importance for feature channels. For the spatial importance, because the softmax attention typically yields sparse attention maps, we propose a loosened attention map for computing the spatial importance, which can better reflect the importance of more regions, as compared to directly using the strict softmax attention. Then our Grad-ECLIP explanation map is calculated with the values in the attention layer as the feature map, weighted by the channel and spatial importances. The same method used to generate the explanation of the image encoder can also be applied to the text encoder to obtain a textual explanation for CLIP. Note that Grad-ECLIP is result-specific and is suitable for both the image and text encoders, i.e., the visual explanation on image is text-specific and the textual explanation of a sentence is image-specific. For the example shown in Fig. 1i, the heat map on the image shows the important region when matching the image with the specific text “a dog is playing with frisbee”, while the degree of green color on the sentence represents the important words, where the most important words “dog” and “frisbee” correspond to the highlighted regions in the image heat map.
@@
Finally, we also present an application of Grad-ECLIP to fine-tuning the CLIP model to boost the fine-grained understanding. Since the ViT-based CLIP model has been shown to have limitations in producing dense representations, due to the pretraining focusing on the whole image-text matching [13, 12, 27, 28], we propose to generate detailed region-phrase corresponding pairs via Grad-ECLIP so as to enhance the fine-grained understanding of CLIP model during fine-tuning. Experiments with zero-shot region classification and down-stream open-vocabulary detection application show that the Grad-ECLIP-enabled fine-tuning is effective.
In summary, the contributions of this paper are five-fold:
@@
    We investigate Grad-ECLIP, a gradient-based visual and textual explanation approach for CLIP to produce high-quality result-specific heat maps for explaining the matching of image-text pairs.
@@
    We show the generalizability of Grad-ECLIP by presenting explanations on datasets of different domains, and verifying the applicability on both ViT and CNN-based CLIP, ViT classifier, and other transformer-based vision language models.
@@
    By using Grad-ECLIP, we explore the properties of CLIP, and reveal the model’s ability of concept decomposition and addibility, strengths and weaknesses in attribution identification, as well as the relationship between word usage and concreteness in the image and text matching.
  5. 5.
    We propose an application of Grad-ECLIP to boost the fine-grained understanding of CLIP via fine-tuning, which adopts the high-quality visual explanation map generated by Grad-ECLIP to produce detailed matching relationships between image regions and the corresponding semantic concepts.
@@
A preliminary version of our work appears in [29]. The extensions over the conference version are as the follows. First, we provide a more detailed derivation of Grad-ECLIP, based on the transformer model. Second, we enrich the qualitative evaluation and verify the generalizability of Grad-ECLIP by adding: (1) the visualization on diverse datasets of different domains; (2) application on ViT-based classifier and other vision language models; (3) adaptation with CNN-based CLIP. Third, we include new ablation studies for the Grad-ECLIP design, including: (1) the effect of the proposed loosen spatial weight; (2) the influence of number of layers involved in the calculation; (3) the influence of multi-attention heads on the visual explanation results. Fourth, we add a new exploration about the relationship between word concreteness and word usage in image-text matching with CLIP. Moreover, we introduce an application of Grad-ECLIP, which applies the visual explanation map to boost the fine-grained region and text matching when fine-tuning CLIP.
@@
The remainder of this paper is organized as follows. The related works about CLIP, explainability in computer vision and fine-grained understanding in CLIP are briefly reviewed in §2. Grad-ECLIP is introduced in §3, and the experiment results are presented in §4, including qualitative evaluations, quantitative evaluations and ablation studies. We then perform analysis of CLIP based on the proposed Grad-ECLIP in §5. Finally, we introduce an application of Grad-ECLIP in §6 for boosting the fine-grained understanding of CLIP.
2 Related Works
@@
We first briefly review the CLIP model, and then discus different types of visual explanation methods in computer vision. Finally, the related works for fine-grained image annotation and understanding in CLIP are introduced.
2.1 Contrastive language-image pre-training
@@
Many multi-modal works have been developed and focus on the interaction of computer vision and natural language processing, such as text-image retrieval [30], image captioning [31], visual question answering [32], and visual grounding [33]. Contrastive language-image pre-training (CLIP) performs contrastive learning on very large-scale web-curated image-text pairs. It shows promising pre-trained representations with superior zero-shot transfer ability on diverse datasets and impressive fine-tuning performance on various downstream tasks. Subsequent works extend and improve CLIP from different aspects: [9, 10] improve the aspects of prompt design and optimization; [7, 8] unifies the vision-language understanding and generation by adding text decoders with image-text cross-attention during pre-training; [11, 12, 13, 28] builds an alignment between region feature or position information with fine-grained object descriptions. Although significant results have been achieved with CLIP and its development, less effort and exploration is focused on its interpretability through visual explanations. In this paper, we propose a novel visual explanation method, which generates high-quality heat maps that reveal the important regions or words used for CLIP’s scoring of an image-text pair.
2.2 Explainability in computer vision
@@
CAM methods, such as CAM [24], Grad-CAM [20], and Grad-CAM++ [34], generate the explanation heat map from a selected feature layer by linearly aggregating the activation maps with weights that indicates each feature’s importance. Grad-CAM computes the weights with global average pooling on the gradients of the model’s prediction w.r.t. the feature layer. Gradient-free CAMs [35, 36, 37] generate weights from the prediction score changes when perturbing features.
Perturbation-based methods [38, 21, 39, 40, 41, 42, 43] perturb the input and observe the changes in output scores to determine the importance of input regions. Such black-box methods are intuitive and highly generalizable to different architecture and tasks, but computationally intensive. The quality of these methods are often greatly influenced by the type or resolution of the perturbations used. While having solid theoretical justification, Shapley-value methods [40] also suffer from large computational complexity.
@@
These previous works are mainly proposed for interpreting CNN-based models. Due to the introduction of self-attention mechanisms in Transformers, recent works [49, 50, 51] have also looked at visual explanations for the Transformer architecture. [14] proposed an Attention flow and Rollout method, which is based on all attention maps in the forward process of model. Since Rollout is class-agnostic, Transformer interpretability [15] and GAME [16] build class-specific relevance-based method for explaining transformer with the internal attention mechanism. However, we found that the explanation methods relying on attention maps in Transformer cannot generate satisfactory results with CLIP, possibly because the sparse attention patterns on the $\mathrm{softmax}$ map. The recent M2IB [22] applies information bottleneck principle to CLIP, which develop an optimization objective to find the compressed representations for both image features and text features. However, a series of hyper-parameters are adopted during the optimization, which limits the generalization in practical applications.
@@
Finally, existing approaches for explaining CLIP [23, 18, 19], which use the cosine similarity map between the image local features and the text features as the explanation map, have the disadvantage that they are only based on the forward (bottom-up) process and thus the attended features are not necessarily used in the final prediction. In contrast, we propose Grad-ECLIP as an effective approach to interpret CLIP, which highlight features that have largest influence on the prediction as measured by the gradient, which is a top-down process.
2.3 Fine-grained image understanding with CLIP
CLIP and its variants [9, 7, 8] exhibit strong representation capabilities and exceptional generalizability through learning general visual-language representations by pre-training on noisy large-scale datasets. Despite the great achievements, CLIP has shown lack of the fine-grained alignment between image regions and text [13, 12, 27, 28] due to its image-level training, which matches an image as a whole to a text description. Thus, the model is unable to generate precise representations of an image region for grounding textual concepts, which will limit the performance of CLIP on the downstream tasks that require region-aware ability. For example, in dense prediction tasks, e.g. object detection and segmentation, the CLIP model is usually utilized as a classifier [52, 53] or the teacher in distillation [54, 55] to process cropped object patches to obtain region features. Some works such as F-vlm [56], CORA [57] and FC-Clip [58] adopt the frozen CLIP model as backbone to produce spatial feature maps, but they all choose the CNN-based CLIP, which can preserve more position information than the vision transformer (ViT-based) architecture. However, due to the image-level training, CLIP models still lack fine-grained alignment and are poor at generating precise image region representations [13, 12, 27, 59].
To mitigate this issue, recent works enhance the fine-grained understanding ability of CLIP by leveraging region-text alignments during pre-training [60, 11, 61, 13, 59]. Since no region-text annotations are provided in the image-text pair training data, most of these methods need to generate image regions with the corresponding text tags using off-the-shelf methods. Some works utilize the annotations in visual grounding datasets [62, 59, 63] or generate pseudo region-text pairs [60, 11, 61] with the help of high-performance detectors that are trained with a large number of object categories. RegionCLIP [13] adopts RPN [64] object proposals while PTP [12] coarsely crops patches, and then they both use CLIP as a classifier to obtain region labels with a large pre-defined pool of concepts, which are parsed from a text corpus. These methods inevitably cost significant extra time and space to preprocess the region annotations, which cannot be neglected when using a huge amount of training data. Moreover, the range of concepts is also limited by the number of pre-defined categories. Another work CLIPSelf [28] facilitates the transfer of the global features of the cropped regions to dense feature extraction by self-distillation, which enhances the local representations during the fine-tuning of CLIP model. However, preprocessing of generating region proposals by a well-trained detector is still required to obtain superior distillation performance – when using randomly cropped patches, the performance significantly drops.
In this paper, we build a novel fine-tuning framework for boosting the fine-grained understanding in CLIP, which adopts the proposed Grad-ECLIP to generate region-aware attention maps for aligning with the corresponding text phrases (concepts). By simply inserting the Grad-ECLIP-based module into the fine-tuning, our method circumvents the resource-consuming preparation of the region annotations and the requirement of high-performance detectors. The inputs of the proposed fine-tuning framework are still image-text pairs, just the same as in the pre-training of CLIP.
@@ 2.3 Fine-grained image understanding with CLIP
3 Grad-ECLIP: Gradient-based visual explanation for CLIP
Our method serves as a gradient-based visual and textual explanation for interpreting the CLIP matching performed on image-text pairs. We start with a brief introduction of CLIP. Then, by decomposing the layers of the transformer and exploring the relationship between the final output and intermediate features, we derive our gradient-based explanation for CLIP (Grad-ECLIP).
3.1 Preliminary on CLIP

### v2

# Grad-ECLIP: Gradient-based Visual and Textual Explanations for CLIP

@@ Abstract
Significant progress has been made in the improvement and downstream applications of the Contrastive Language-Image Pre-training (CLIP) vision-language model, while less attention has been paid to the interpretation of CLIP. We propose a Gradient-based visual and textual Explanation method for CLIP (Grad-ECLIP), which interprets the matching result of CLIP for a specific input image-text pair. By decomposing the encoder’s architecture and identifying the relationship between matching similarity and intermediate spatial features, Grad-ECLIP generates effective heat maps that reveal the impact of image regions or words on the CLIP results. Unlike previous Transformer interpretation methods that focus on utilizing self-attention maps, which are typically extremely sparse in CLIP, we produce high-quality visual explanations by applying channel and spatial weights to token features. Qualitative and quantitative evaluations verify the effectiveness and superiority of Grad-ECLIP compared with the state-of-the-art methods. Finally, a series of analyses are conducted based on our visual and textual explanation results, from which we explore the working mechanism of image-text matching, the strengths and limitations in attribution identification of CLIP, and the relationship between the concreteness/abstractness of a word and its usage in CLIP. The code of Grad-ECLIP is available here: https://github.com/Cyang-Zhao/Grad-Eclip.
I Introduction
@@
Recently, Vision-Language Pre-training (VLP) models have bridged the gap between the low-level visual information in pixels and high-level conceptual knowledge embedded in language. By leveraging self-supervised objectives on web-scale data, VLP models learn an aligned understanding of both vision and language, and can flexibly perform applications on various tasks, such as classification, image-text retrieval, visual questioning answering (VQA), image captioning, and serve as foundational technology for the visual inputs to generative AI systems. The evolution of VLP has been marked by several influential architectures. The Contrastive Language-Image Pre-training (CLIP) [40] model, which employs a dual-encoder architecture trained with a contrastive objective to facilitate image-text matching, is a notable milestone. CLIP significantly improves performance on various downstream tasks [7, 6, 33, 59, 64], utilizing both zero-shot and fine-tuning methodologies. Inspired by CLIP, VLP has been further developed by exploring different perspectives, including unifying vision-language understanding and generation [66, 26], prompt design [74, 11], and region-aware enhancement [28, 56, 72]. Although researchers devote many efforts into improving multi-modal pre-training or exploring the usages in downstream tasks, less attention has been focused on the interpretation or explanation of the learned vision and language matching process.
@@
Figure 1: Visual and textual explanations of CLIP for the image with the text “A dog is playing with frisbee” using (a) CLIPSurgery [29]; (b) MaskCLIP [73]; (c) Grad-CAM [46]; (d) RISE [36]; (e) raw attention in the last layer; (f) Rollout [1]; (g) GAME [9]; (h) M2IB [58]; and (i) Our Grad-ECLIP. For (e) to (i), textual explanations on the sentence are shown, where the degree of green color represents the word importance. Other methods (a-d) are not applicable to text. The faithfulness (IMD) of each method on the image is shown under the heat map of each method (higher is better).
Previous visual explanation works have considered interpreting the transformer architecture used by VLP models. Attention Rollout [1] generates explanations by aggregating attention maps computed along the forward pass of the model. Relevance-based methods [10, 9] apply Layer-wise Relevance Propagation (LRP) [3] and also rely on the attention mechanism in the model architecture. Since Rollout and many LRP variants are class-agnostic, Transformer interpretability [10] and Generic Attention-Model Explainability (GAME) [9] build class-specific relevance-based explanations using the self-attention or co-attention. However, just treating VLP models as a vision transformer and generating explanations based on self-attention sometimes leads to low-faithfulness and confusing visualization results because of sparse attention maps (see Fig. 1e-g). Figure 1 shows the visual explanation heat maps on CLIP with the evaluated IMD faithfulness, where the value represents an average influence on the image-text similarity score within 100 steps when deleting and inserting image pixels based on the heat maps (see §IV-B1 for more details about the IMD metric). Higher IMD values are better, indicating better fidelity to the actual important regions.
@@
The vision-language matching interpretation is explored by ECLIP [30] and CLIPSurgery [29] with CLIP by computing an image-text similarity map. They attempt to solve the counter-intuitive problem that background patch features get higher similarity with the text feature than the foreground. However, to obtain reasonable similarity maps, new additional projection layers or changing the structure of the original CLIP are required. Although the parameters of CLIP encoders are frozen, learning more black-box parameters with extra data or modifying the original model architecture makes the explanation less interpretable. MaskCLIP [73] also provides a technique to calculate class-specific image-text similarity map by passing the value features of the last attention layer through later linear layers as image patch features. Although the similarity maps from CLIPSurgery and MaskCLIP are able to localize the concept in the text (see Fig. 1a-b), lower faithfulness results come from a noisy background and confusingly highlight points on the location unrelated to the explained target matching text. The disadvantage of these similarity-map methods is that they are only forward processing, and the attended features are not necessarily used in the final prediction.
@@
To better focus on the discriminative features used in the prediction, gradient-based methods with class-activation maps (CAM) [68], such as Grad-CAM [46], Layer-CAM [21], and FullGrad [49], consider the gradient of the prediction with respect to features from a CNN layer as weights, and locates the class-specific discriminative regions by weighted aggregation of the feature maps. Fig.1c shows the faithfulness and visualization when adapting Grad-CAM on CLIP, where the cosine similarity of the image-text pair is adopted as the prediction and the gradients are calculated w.r.t. the patch tokens from the ViT layers. Since there are no gradients w.r.t. the patch tokens in the final layer because they are not involved in the calculation of the matching score, feature outputs from the penultimate layer of ViT are adopted. The results of Grad-CAM do not well explain CLIP, and suffer from the same problem of highlighting unrelated points as MaskCLIP, suggesting that the layer features of ViT are unsuitable for CAM methods.
@@
In this paper, we explore a more effective way to interpret the image-text matching in VLP models. Using the representative CLIP dual-encoders as the paradigm architecture, we analyze how CLIP obtains the final feature embedding and derive the relationship between the image and text embedding similarity score and intermediate features via a series of approximations. We first make the derivation based on the more widely used ViT-based CLIP, which is suitable for both the transformer-based image and text encoders. Then, we give the adaptation of the CNN-based architecture. Based on the CAM principle, we then propose a novel gradient-based visual explanation method for CLIP (Grad-ECLIP), which generates the importance heat map by aggregating the intermediate image features with result-related channel and spatial importance weights. Our proposed method uses the gradients of the image-text matching score w.r.t. the attention layer as the importance for feature channels. For the spatial importance, because the softmax attention typically yields sparse attention maps, we propose a loosened attention map for computing the spatial importance, which can better reflect the importance of more regions, as compared to directly using the strict softmax attention. Then our Grad-ECLIP explanation map is calculated with the values in the attention layer as the feature map, weighted by the channel and spatial importances. The same method used to generate the explanation of the image encoder can also be applied to the text encoder to obtain a textual explanation for CLIP. Note that Grad-ECLIP is result-specific and is suitable for both the image and text encoders, i.e., the visual explanation on image is text-specific and the textual explanation of a sentence is image-specific. As the example shows in Fig. 1i, Grad-ECLIP produces the best faithfulness result on this example image-text pair. The heat map on the image shows the important region when matching the image with the specific text “a dog is playing with frisbee”, while the degree of green color on the sentence represents the important words, where the most important words “dog” and “frisbee” correspond to the highlighted regions in the image heat map.
@@
Finally, we also present an application of Grad-ECLIP to fine-tuning the CLIP model to boost the fine-grained understanding. Since the ViT-based CLIP model has been shown to have limitations in producing dense representations, due to the pretraining focusing on the whole image-text matching [72, 56, 23, 62], we propose to generate detailed region-phrase corresponding pairs via Grad-ECLIP so as to enhance the fine-grained understanding of the CLIP model during fine-tuning. Experiments with zero-shot region classification and downstream open-vocabulary detection application show that the Grad-ECLIP-enabled fine-tuning is effective.
In summary, the contributions of this paper are fivefold:
@@
    We explore the explanation of image-text matching in VLP models by investigating Grad-ECLIP, a gradient-based visual and textual explanation approach for CLIP dual-encoders to produce high-quality result-specific heat maps for explaining the matching of image-text pairs.
@@
    We show the generalizability of Grad-ECLIP by verifying the applicability on both ViT and CNN-based CLIP, other transformer-based VLP models like BLIP, as well as ViT classifier, and presenting explanations on datasets of different domains.
@@
    By using Grad-ECLIP, we explore the properties of CLIP, and reveal the model’s ability of concept decomposition and addability, strengths and weaknesses in attribution identification, as well as the relationship between word usage and concreteness in the image and text matching.
@@
A preliminary version of our work appears in [71]. The extensions over the conference version are as follows. First, we provide a more detailed derivation of Grad-ECLIP, based on the transformer model. Second, we enrich the qualitative evaluation and verify the generalizability of Grad-ECLIP by adding: (1) the visualization on diverse datasets of different domains; (2) application on ViT-based classifier and other vision language models; (3) adaptation with CNN-based CLIP. Third, we include new ablation studies for the Grad-ECLIP design, including: (1) the effect of the proposed loosen spatial weight; (2) the influence of the number of layers involved in the calculation; (3) the influence of multi-attention heads on the visual explanation results. Fourth, we add quantitative experiments for the analysis of CLIP and a new exploration about the relationship between word concreteness and word usage in image-text matching with CLIP. Moreover, we conduct a user study to evaluate the trustability of the visual explanations.
@@
The remainder of this paper is organized as follows. The related works about CLIP and explainability in computer vision are briefly reviewed in §II. Grad-ECLIP is introduced in §III, and the experiment results are presented in §IV, including qualitative evaluations, quantitative evaluations, and ablation studies. Finally, we perform analysis of CLIP based on the proposed Grad-ECLIP in §V.
II Related Works
@@
We first briefly review the CLIP model, and then discus different types of visual explanation methods in computer vision.
II-A Contrastive language-image pre-training
@@
Many multi-modal works have been developed and focus on the interaction of computer vision and natural language processing, such as text-image retrieval [61], image captioning [65], visual question answering [2], and visual grounding [38]. Contrastive language-image pre-training (CLIP) performs contrastive learning on very large-scale web-curated image-text pairs. It shows promising pre-trained representations with superior zero-shot transfer ability on diverse datasets and impressive fine-tuning performance on various downstream tasks. Subsequent works extend and improve CLIP from different aspects: [74, 11] improve the aspects of prompt design and optimization; [66, 26] unify the vision-language understanding and generation by adding text decoders with image-text cross-attention during pre-training; [28, 56, 72, 62] build an alignment between region feature or position information with fine-grained object descriptions. Although significant results have been achieved with CLIP and its development, less effort and exploration are focused on its interpretability through visual explanations. In this paper, we propose a novel visual explanation method, which generates high-quality heat maps that reveal the important regions or words used for CLIP’s scoring of an image-text pair.
II-B Explainability in computer vision
@@
CAM methods, such as CAM [68], Grad-CAM [46], and Grad-CAM++ [8], generate the explanation heat map from a selected feature layer by linearly aggregating the activation maps with weights that indicate each feature’s importance. Grad-CAM computes the weights with global average pooling on the gradients of the model’s prediction w.r.t the feature layer. Gradient-free CAMs [41, 54, 53] generate weights from the prediction score changes when perturbing features.
Perturbation-based methods [42, 36, 13, 32, 52, 25, 37] perturb the input and observe the changes in output scores to determine the importance of input regions. Such black-box methods are intuitive and highly generalizable to different architectures and tasks, but computationally intensive. The quality of these methods is often greatly influenced by the type or resolution of the perturbations used. While having solid theoretical justification, Shapley-value methods [32] also suffer from large computational complexity.
@@
These previous works are mainly proposed for interpreting CNN-based models. Due to the introduction of self-attention mechanisms in Transformers, recent works [39, 63, 67] have also looked at visual explanations for the Transformer architecture. [1] proposed an Attention flow and Rollout method, which is based on all attention maps in the forward process of the model. Since Rollout is class-agnostic, Transformer interpretability [10] and GAME [9] build class-specific relevance-based methods for explaining the transformer with the internal attention mechanism. However, we found that the explanation methods relying on attention maps in Transformer cannot generate satisfactory results with CLIP, possibly because of the sparse attention patterns on the $\mathrm{softmax}$ map. The recent M2IB [58] applies the information bottleneck principle to CLIP, which develops an optimization objective to find the compressed representations for both image features and text features. However, a series of hyperparameters are adopted during the optimization, which limits the generalization in practical applications.
@@
Finally, existing approaches for explaining CLIP [30, 29, 73], which use the cosine similarity map between the image local features and the text features as the explanation map, have the disadvantage that they are only based on the forward (bottom-up) process and thus the attended features are not necessarily used in the final prediction. In contrast, we propose Grad-ECLIP as an effective approach to interpret CLIP, which highlights features that have the largest influence on the prediction as measured by the gradient, which is a top-down process.
@@
III Grad-ECLIP: Gradient-based visual explanation for CLIP
Our method serves as a gradient-based visual and textual explanation for interpreting the CLIP matching performed on image-text pairs. We start with a brief introduction of preliminaries. Then, by decomposing the layers of the transformer and exploring the relationship between the final output and intermediate features, we derive our gradient-based explanation for CLIP (Grad-ECLIP).
III-A Preliminaries